OpenStudy (anonymous):

Give the answer with the appropriate unit of measurement. I'll post the equation below @ags2658

OpenStudy (anonymous):

\[(\frac{ 9 feet }{ 2 seconds } ) (\frac{ 3,600 seconds }{ 1 hour })(\frac{ 1mile }{ 5,280 feet })\]

OpenStudy (anonymous):

Okay. Feet cancel out in this equation, since they are BOTH in the numerator and the denominator. The same thing is true for seconds. Therefore, our final unit is MILES per HOUR. I don't have time to do the math, since I am leaving now. However, I trust that you can do it. The top is 32400 and the bottom is 10560. See ya later! :D

OpenStudy (anonymous):

Wait, with the little time that I have left, I did the division for you. It is approximately 3 miles per hour. To be more precise, it is 3.068181818.... :D